# Boundary phase transitions of two-dimensional quantum critical XXZ model

### v1

@@ 3.1 Two-leg ladder
Introducing the symmetric and antisymmetric combinations of boson fields, $\phi_{\mathrm{S,A}}=\frac{1}{\sqrt{2}}(\phi_{1}\pm\phi_{2})$ and $\theta_{\mathrm{S,A}}=\frac{1}{\sqrt{2}}(\theta_{1}\pm\theta_{2})$, and retaining only the most relevant potential terms, we find the following bosonized Hamiltonian of the two-leg ladder in the continuum limit [35, 36, 37, 38, 39],
$\begin{split}H_{0}=&\int\mathrm{d}x\,\Bigg{(}\sum_{i=\mathrm{S,A}}\Big{(}\frac{1}{2K_{i}^{2}}(\partial_{x}\phi_{i})^{2}+\frac{1}{2}(\partial_{x}\theta_{i})^{2}\Big{)}+2g_{1}\cos\big{(}\sqrt{2\pi}\theta_{\mathrm{A}}\big{)}\\ &+2g_{2}\cos\big{(}2\sqrt{2\pi}\phi_{\mathrm{A}}\big{)}+2g_{3}\cos\big{(}2\sqrt{2\pi}\phi_{\mathrm{S}}\big{)}\Bigg{)},\end{split}$ (49)
in which the coupling constants are defined by $K_{\mathrm{S,A}}=\big{(}1+\frac{4\Delta}{\pi}\pm\frac{\lambda_{z}}{2\pi}\big{)}^{-1/2}$, $g_{1}=\pi\lambda_{xy}/(2\pi a_{0})^{2}$, and $g_{2}=g_{3}=\lambda_{z}/(2\pi a_{0})^{2}$. The symmetric and the antisymmetric sectors decouple from each other due to the interchain reflection symmetry. The boson fields in each sector satisfy the commutation relations,
@@
We note that the action of the ladder is not Lorentz invariant in general, because the symmetric and the antisymmetric sectors have different velocities, thus we use different definitions of complex coordinates in the two sectors implicitly above. In Sec. 4.2, we will show that the velocity difference diminishes in the RG flow. The correlation functions at the Gaussian fixed point are given by
@@
and the RG flow is plotted in Fig. 3. In region I, $g_{3}$ is relevant and flows to infinity, while $K_{\mathrm{S}}$ vanishes in the RG flow, thus $\phi_{\mathrm{S}}$ is pinned at the minima of the cosine potential, $\phi_{\mathrm{S}}=\sqrt{\pi/2}(n+1/2)$ ($n\in\mathbb{Z}$), while correlation functions of vertex operators of the dual field $\theta_{\mathrm{S}}$ decay exponentially. In region II, $g_{3}$ vanishes and $K_{\mathrm{S}}$ flows to a finite value larger than one, thus the boson fields $\phi_{\mathrm{S}}$ and $\theta_{\mathrm{S}}$ remain critical.
@@
We find at least one of the cosine terms is relevant and pins the corresponding boson field $\phi_{\mathrm{A}}$ or $\theta_{\mathrm{A}}$ at the minima of the potential term. In the first case, $g_{2}$ flows to infinity, thus $\phi_{\mathrm{A}}$ is pinned at $\sqrt{\pi/2}(n+1/2)$ ($n\in\mathbb{Z}$), while $K_{\mathrm{A}}$ and $g_{1}$ vanish in the RG flow. In the second case, $g_{1}$ flows to infinity, thus $\theta_{\mathrm{A}}$ is pinned at $\sqrt{2\pi}(n+1/2)$ ($n\in\mathbb{Z}$), while $K_{\mathrm{A}}$ flows to infinity and $g_{2}$ vanishes.
Combining the results of the symmetric and the antisymmetric sectors, we find four phases of the two-leg ladder. In order to characterize these phases, we define $\vec{S}=\vec{\sigma}-\vec{\tau}$ and express it in terms of boson fields,
$\begin{split}S^{+}(x)=&\frac{2i}{\sqrt{2\pi\alpha}}e^{i\sqrt{\pi/2}\theta_{\mathrm{S}}(x)}\bigg{(}e^{i\sqrt{2\pi}\phi_{\mathrm{S}}(x)}\sin\Big{(}\sqrt{\pi/2}\theta_{\mathrm{A}}(x)+\sqrt{2\pi}\phi_{\mathrm{A}}(x)\Big{)}\\ +&(-1)^{x/a_{0}}\sin\Big{(}\sqrt{\pi/2}\theta_{\mathrm{A}}(x)\Big{)}\bigg{)},\end{split}$ (61)
@@
Ising-AF phase. (Symmetric sector: $g_{3}\rightarrow+\infty$, $K_{\mathrm{S}}\rightarrow 0$; Antisymmetric sector: $g_{1}\rightarrow 0$, $g_{2}\rightarrow+\infty$, $K_{\mathrm{A}}\rightarrow 0$.) The boson fields $\phi_{\mathrm{S}}$ and $\phi_{\mathrm{A}}$ are pinned at the minima of their cosine potential terms, and the correlation functions of the vertex operators of the dual fields $\theta_{\mathrm{S}}$ and $\theta_{\mathrm{A}}$ decay exponentially. Therefore, there is a long-range easy-axis AF order, $\langle(-1)^{x/a_{0}}S^{z}(x)\rangle\neq 0$, and the inplane spin correlation function $\langle S^{+}S^{-}+S^{-}S^{+}\rangle$ decays exponentially.
Singlet phase. ($g_{3}\rightarrow+\infty$, $K_{\mathrm{S}}\rightarrow 0$; $g_{1}\rightarrow+\infty$, $g_{2}\rightarrow 0$, $K_{\mathrm{A}}\rightarrow+\infty$.) The boson fields $\phi_{\mathrm{S}}$ and $\theta_{\mathrm{A}}$ are pinned at the minima of their corresponding cosine potential terms, and the correlations of the vertex operators of their dual fields $\theta_{\mathrm{S}}$ and $\phi_{\mathrm{A}}$ decay exponentially. Therefore, we find both spin correlation functions $\langle S^{z}S^{z}\rangle$ and $\langle S^{+}S^{-}+S^{-}S^{+}\rangle$ decay exponentially, and this phase is a disordered phase.
XY1 phase. ($g_{3}\rightarrow 0$, $K_{\mathrm{S}}\rightarrow\text{finite value}$; $g_{1}\rightarrow+\infty$, $g_{2}\rightarrow 0$, $K_{\mathrm{A}}\rightarrow+\infty$.) The boson field $\theta_{\mathrm{A}}$ is pinned, while $\phi_{\mathrm{S}}$ and $\theta_{\mathrm{S}}$ remain critical. Therefore, $\langle S^{z}S^{z}\rangle$ decays exponentially, while the AF component of the inplane spin correlation function $\langle S^{+}S^{-}+S^{-}S^{+}\rangle$ decays in power law.
XY2 phase. ($g_{3}\rightarrow 0$, $K_{\mathrm{S}}\rightarrow\text{finite value}$; $g_{1}\rightarrow 0$, $g_{2}\rightarrow+\infty$, $K_{\mathrm{A}}\rightarrow 0$.) The boson field $\phi_{\mathrm{A}}$ is pinned, while $\phi_{\mathrm{S}}$ and $\theta_{\mathrm{S}}$ remain critical. Therefore, $\langle S^{+}S^{-}+S^{-}S^{+}\rangle$ decays exponentially, while the AF component of $\langle S^{z}S^{z}\rangle$ and $\big{\langle}(S^{+})^{2}(S^{-})^{2}+(S^{-})^{2}(S^{+})^{2}\big{\rangle}$ decay in power law.
3.2 Coupling to the bulk
@@
Here, the effective coupling of the lower spin chain to the bulk is antiferromagnetic with $g_{xx}^{\sigma}>0$ and $g_{z}^{\sigma}>0$, while that of the upper spin chain is ferromagnetic with $g_{xx}^{\tau}<0$ and $g_{z}^{\tau}<0$. We will neglect the velocity difference between the two spin chains and the bulk for the moment and turn to this issue in Sec. 4.2.
The RG equations can be derived from the OPE relations listed in Appendix B.2. Defining $g_{i}^{\pm}=g_{i}^{\sigma}\pm g_{i}^{\tau}$ with $i=xx,z$, we find
@@
Let us examine the stability of the four phases of the two-leg ladder in turn.
Singlet phase. In this phase, $K_{\mathrm{S}}\rightarrow 0$, $K_{\mathrm{A}}\rightarrow+\infty$, $g_{1}\rightarrow 0$, $g_{2}\rightarrow+\infty$, and $g_{3}\rightarrow+\infty$ in the ladder. It is easy to see that $g_{xx}^{+}$ and $g_{z}^{+}$ are irrelevant and vanish in the RG flow. In order to find out the fate of other coupling constants, we take a closer look at the asymptotic form of the RG equations,
@@
thus we find $g_{1}\sim e^{2l}$ while $K_{\mathrm{S}}^{-1}\sim e^{4l}$ diverges even faster, and $g_{xx}^{-}$ vanishes in the RG flow. With similar analysis, we find $g_{z}^{-}$ vanishes as well. Therefore, all coupling terms between the ladder and the bulk are irrelevant, and the boundary retains the ordinary critical behavior without breaking any symmetry.
Ising-AF phase. In this phase, $K_{\mathrm{S}}\rightarrow 0$, $K_{\mathrm{A}}\rightarrow 0$, $g_{1}\rightarrow 0$, $g_{2}\rightarrow+\infty$, and $g_{3}\rightarrow+\infty$ in the ladder. It is easy to see that $g_{xx}^{\pm}$ and $g_{z}^{+}$ vanish, while $g_{z}^{-}$ increases in the RG flow, which turns out to stabilize the original RG flow of the coupling terms in the ladder. Therefore, the coupling terms $g_{xx}^{\pm}$ are irrelevant and the inplane spin correlation on the boundary remains in the ordinary universality class, while it develops long-range easy-axis AF order with an extraordinary critical behavior.
@@
The coupling constants $K_{\mathrm{S,A}}$, $g_{1}$ and $g_{xx}^{-}$ diverge at a finite renormalization scale $l_{0}<+\infty$ implying the inadequacy of the perturbative RG analysis, thus the Callan-Symanzik equation cannot be integrated to find the asymptotic behavior of the correlation functions. Nonetheless, the strong divergence of $K_{\mathrm{S}}$ and $K_{\mathrm{A}}$ suggest that both dual fields $\theta_{\mathrm{S}}$ and $\theta_{\mathrm{A}}$ are pinned at static values, thus the boundary has a long-range easy-plane AF order, and shows extraordinary critical behavior.
4 Irrelevance of weak velocity difference

### v3

@@ III.1 Two-leg ladder
Introducing the symmetric and antisymmetric combinations of boson fields, $\phi_{\mathrm{S,A}}=\frac{1}{\sqrt{2}}(\phi_{1}\pm\phi_{2})$ and $\theta_{\mathrm{S,A}}=\frac{1}{\sqrt{2}}(\theta_{1}\pm\theta_{2})$, and retaining only the most relevant potential terms, the following bosonized Hamiltonian of the two-leg ladder in the continuum limit is obtained [35, 36, 37, 38, 39, 40, 41, 42],
$H_{0}=\int\mathrm{d}x\,\Bigg{(}\sum_{i=\mathrm{S,A}}\Big{(}\frac{1}{2K_{i}^{2}}(\partial_{x}\phi_{i})^{2}+\frac{1}{2}(\partial_{x}\theta_{i})^{2}\Big{)}+2g_{1}\cos\big{(}\sqrt{2\pi}\theta_{\mathrm{A}}\big{)}+2g_{2}\cos\big{(}2\sqrt{2\pi}\phi_{\mathrm{A}}\big{)}+2g_{3}\cos\big{(}2\sqrt{2\pi}\phi_{\mathrm{S}}\big{)}\Bigg{)},$ (50)
in which the coupling constants are defined by $K_{\mathrm{S,A}}=\big{(}1+\frac{4\Delta}{\pi}\pm\frac{\lambda_{z}}{2\pi}\big{)}^{-1/2}$, $g_{1}=\pi\lambda_{xy}/(2\pi a_{0})^{2}$, and $g_{2}=g_{3}=\lambda_{z}/(2\pi a_{0})^{2}$. The $g_{a}$ ($1\leq a\leq 3$) are positive for AF interchain coupling and negative for ferromagnetic (FM) interchain coupling. The symmetric and the antisymmetric sectors decouple from each other due to the interchain reflection symmetry. The boson fields in each sector satisfy the commutation relations,
@@
We note that the action of the ladder is not Lorentz invariant in general, because the symmetric and the antisymmetric sectors can have different velocities, thus we use different definitions of complex coordinates in the two sectors implicitly. In Sec. IV.2, we will show that the velocity difference diminishes in the RG flow. The correlation functions at the Gaussian fixed point are given by
@@
[…]
@@
[…]
Combining the results of the symmetric and the antisymmetric sectors, we find four phases of the two-leg ladder. In order to characterize these phases, we define $\vec{S}=\vec{\sigma}-\vec{\tau}$ if the interchain coupling is antiferromagnetic and express it in terms of boson fields,
$S^{+}(x)=\frac{2i}{\sqrt{2\pi\alpha}}e^{i\sqrt{\pi/2}\theta_{\mathrm{S}}(x)}\bigg{(}e^{i\sqrt{2\pi}\phi_{\mathrm{S}}(x)}\sin\Big{(}\sqrt{\pi/2}\theta_{\mathrm{A}}(x)+\sqrt{2\pi}\phi_{\mathrm{A}}(x)\Big{)}+(-1)^{x/a_{0}}\sin\Big{(}\sqrt{\pi/2}\theta_{\mathrm{A}}(x)\Big{)}\bigg{)},$ (62)
@@
In the case of FM interchain coupling, we define $\vec{S}=\vec{\sigma}+\vec{\tau}$, and
$S^{+}(x)=\frac{2}{\sqrt{2\pi\alpha}}e^{i\sqrt{\pi/2}\theta_{\mathrm{S}}(x)}\bigg{(}e^{i\sqrt{2\pi}\phi_{\mathrm{S}}(x)}\cos\Big{(}\sqrt{\pi/2}\theta_{\mathrm{A}}(x)+\sqrt{2\pi}\phi_{\mathrm{A}}(x)\Big{)}+(-1)^{x/a_{0}}\cos\Big{(}\sqrt{\pi/2}\theta_{\mathrm{A}}(x)\Big{)}\bigg{)},$ (64)
$S^{z}(x)=\sqrt{\frac{8}{\pi}}\partial_{x}\phi_{\mathrm{S}}+(-1)^{x/a_{0}}\frac{4}{\pi\alpha}\cos\big{(}\sqrt{2\pi}\phi_{\mathrm{S}}\big{)}\cos\big{(}\sqrt{2\pi}\phi_{\mathrm{A}}\big{)}.$ (65)
Ising-AF phase. (Symmetric sector: $g_{3}\rightarrow\pm\infty$, $K_{\mathrm{S}}\rightarrow 0$; Antisymmetric sector: $g_{1}\rightarrow 0$, $g_{2}\rightarrow\pm\infty$, $K_{\mathrm{A}}\rightarrow 0$.) The boson fields $\phi_{\mathrm{S}}$ and $\phi_{\mathrm{A}}$ are pinned at the minima of their cosine potential terms, and the correlation functions of the vertex operators of the dual fields $\theta_{\mathrm{S}}$ and $\theta_{\mathrm{A}}$ decay exponentially. Therefore, there is a long-range easy-axis AF order, $\langle(-1)^{x/a_{0}}S^{z}(x)\rangle\neq 0$, and the inplane spin correlation function $\langle S^{+}S^{-}+S^{-}S^{+}\rangle$ decays exponentially.
Singlet phase. ($g_{3}\rightarrow\pm\infty$, $K_{\mathrm{S}}\rightarrow 0$; $g_{1}\rightarrow\pm\infty$, $g_{2}\rightarrow 0$, $K_{\mathrm{A}}\rightarrow+\infty$.) The boson fields $\phi_{\mathrm{S}}$ and $\theta_{\mathrm{A}}$ are pinned at the minima of their corresponding cosine potential terms, and the correlations of the vertex operators of their dual fields $\theta_{\mathrm{S}}$ and $\phi_{\mathrm{A}}$ decay exponentially. Therefore, we find both spin correlation functions $\langle S^{z}S^{z}\rangle$ and $\langle S^{+}S^{-}+S^{-}S^{+}\rangle$ decay exponentially, and this phase is a disordered phase. This phase with a FM interchain coupling smoothly connects to the Haldane phase of the spin-$1$ Heisenberg chain [43, 44].
XY1 phase. ($g_{3}\rightarrow 0$, $K_{\mathrm{S}}\rightarrow\text{finite value}$; $g_{1}\rightarrow\pm\infty$, $g_{2}\rightarrow 0$, $K_{\mathrm{A}}\rightarrow+\infty$.) The boson field $\theta_{\mathrm{A}}$ is pinned, while $\phi_{\mathrm{S}}$ and $\theta_{\mathrm{S}}$ remain critical. Therefore, $\langle S^{z}S^{z}\rangle$ decays exponentially, while the AF component of the inplane spin correlation function $\langle S^{+}S^{-}+S^{-}S^{+}\rangle$ decays in power law.
XY2 phase. ($g_{3}\rightarrow 0$, $K_{\mathrm{S}}\rightarrow\text{finite value}$; $g_{1}\rightarrow 0$, $g_{2}\rightarrow\pm\infty$, $K_{\mathrm{A}}\rightarrow 0$.) The boson field $\phi_{\mathrm{A}}$ is pinned, while $\phi_{\mathrm{S}}$ and $\theta_{\mathrm{S}}$ remain critical. Therefore, $\langle S^{+}S^{-}+S^{-}S^{+}\rangle$ decays exponentially, while the AF component of $\langle S^{z}S^{z}\rangle$ and $\big{\langle}(S^{+})^{2}(S^{-})^{2}+(S^{-})^{2}(S^{+})^{2}\big{\rangle}$ decay in power law.
III.2 Coupling to the bulk
@@
Here, the effective coupling of the lower spin chain to the bulk is antiferromagnetic with $g_{xx}^{\sigma}>0$ and $g_{z}^{\sigma}>0$. The effective coupling of the upper spin chain to the bulk depends on the interchain coupling: $g_{xx}^{\tau},~{}g_{z}^{\tau}<0$ for AF interchain coupling $g_{a}>0$ ($1\leq a\leq 3$), while $g_{xx}^{\tau},~{}g_{z}^{\tau}>0$ for FM interchain coupling $g_{a}<0$ ($1\leq a\leq 3$). We will neglect the velocity difference between the two spin chains and the bulk for the moment and turn to this issue in Sec. IV.2. The RG equations can be derived from the OPE relations listed in Appendix B.2. Defining $g_{i}^{\pm}=g_{i}^{\sigma}\pm g_{i}^{\tau}$ with $i=xx,z$, we find
@@
Table 2: Summary of possible boundary phases of the 2D quantum critical AF XXZ model weakly coupled with a two-leg ladder on the boundary.
easy-plane AF spin rotation O(2) ordinary extraordinary
singlet none ordinary ordinary
Let us examine the stability of the four phases of the two-leg ladder in turn. The results are summarized in Table 2.
Singlet phase. In this phase, $K_{\mathrm{S}}\rightarrow 0$, $K_{\mathrm{A}}\rightarrow+\infty$, $g_{1}\rightarrow\pm\infty$, $g_{2}\rightarrow 0$, and $g_{3}\rightarrow\pm\infty$ in the ladder. Let us first examine the AF interchain coupling case with $g_{a}>0$ ($1\leq a\leq 3$). It is easy to see that $g_{xx}^{+}$ and $g_{z}^{+}$ are irrelevant and vanish in the RG flow. In order to find out the fate of other coupling constants, we take a closer look at the asymptotic form of the RG equations,
@@
thus we find $g_{1,3}\sim e^{2l}$ while $K_{\mathrm{S}}^{-1}\sim e^{4l}$ diverges even faster, and $g_{xx}^{-}$ vanishes in the RG flow. With similar analysis, we find $g_{z}^{-}$ vanishes as well. In the FM interchain coupling case with $g_{a}<0$ ($1\leq a\leq 3$), the roles of $g_{xx,z}^{+}$ and $g_{xx,z}^{-}$ are interchanged; Nonetheless, the coupling constants $g_{xx}^{\pm}$ and $g_{z}^{\pm}$ vanish with the RG flow altogether. Therefore, all coupling terms between the ladder and the bulk are irrelevant, and the boundary retains the ordinary critical behavior without breaking any symmetry.
Because flipping the signs of $g_{a}$ ($1\leq a\leq 3$) altogether amounts to interchanging the roles of $g_{xx,z}^{+}$ and $g_{xx,z}^{-}$ in the analysis of the coupled RG equations, we find the FM interchain coupling does not affect the main conclusions in each phase discussed below.
Ising-AF phase. In this phase, $K_{\mathrm{S}}\rightarrow 0$, $K_{\mathrm{A}}\rightarrow 0$, $g_{1}\rightarrow 0$, $g_{2}\rightarrow+\infty$, and $g_{3}\rightarrow+\infty$ in the ladder for $g_{a}>0$ ($1\leq a\leq 3$). It is easy to see that $g_{xx}^{\pm}$ and $g_{z}^{+}$ vanish, while $g_{z}^{-}$ increases in the RG flow, which turns out to stabilize the original RG flow of the coupling terms in the ladder. Therefore, the coupling terms $g_{xx}^{\pm}$ are irrelevant and the inplane spin correlation on the boundary remains in the ordinary universality class, while it develops long-range easy-axis AF order with an extraordinary critical behavior.
@@
The coupling constants $K_{\mathrm{S,A}}$, $g_{1}$ and $g_{xx}^{-}$ diverge at a finite renormalization scale $l_{0}<+\infty$ implying the inadequacy of the perturbative RG analysis, thus the Callan-Symanzik equation cannot be integrated to find the asymptotic behavior of the correlation functions. Nonetheless, the strong divergence of $K_{\mathrm{S}}$ and $K_{\mathrm{A}}$ suggests that both dual fields $\theta_{\mathrm{S}}$ and $\theta_{\mathrm{A}}$ are pinned at static values, thereby the boundary has a long-range easy-plane AF order, and shows extraordinary critical behavior.
Figure 4: Numerical solutions of the coupled RG equations (68)–(76). The bare coupling constants are specified as the initial conditions of the equations. Weak couplings to the bulk are introduced into each phase of the dangling ladder. (a, b) Singlet phase. The bare coupling constants of the ladder: $K_{\mathrm{S}}(0)=2$, $K_{\mathrm{A}}=5$, $g_{1}=5$, $g_{2}=0.5$, and $g_{3}=4$. (c, d) Ising-AF phase. The bare coupling constants of the ladder: $K_{\mathrm{S}}(0)=2$, $K_{\mathrm{A}}=1$, $g_{1}=0.5$, $g_{2}=2$, and $g_{3}=4$. (e, f) XY2 phase. The bare coupling constants of the ladder: $K_{\mathrm{S}}(0)=2$, $K_{\mathrm{A}}=1$, $g_{1}=0.5$, $g_{2}=2$, and $g_{3}=0.25$. (g, h) XY1 phase. The bare coupling constants of the ladder: $K_{\mathrm{S}}(0)=2$, $K_{\mathrm{A}}=5$, $g_{1}=5$, $g_{2}=0.5$, and $g_{3}=0.25$. In each phase, the following bare coupling strengths to the bulk are introduced: $g_{xx}^{+}=0.01$, $g_{xx}^{-}=0.01$, $g_{z}^{+}=0.02$, and $g_{z}^{-}=0.02$. The RG flow of the coupling constants confirm the perturbative analysis in the main text.
Figure 5: Numerical solutions of the coupled RG equations (68)–(76). The bare coupling constants are specified as the initial conditions of the equations. A relatively strong coupling to the bulk $g_{xx}^{-}=0.05$ is introduced into the singlet phase of the dangling ladder. Other bare coupling constants are the same as Fig. 4 (a, b). The RG flow of the coupling constants indicates that the singlet phase is destabilized and gives way to an AF ordered state on the boundary.
The coupled RG equations (68)–(76) can be solved numerically with given bare coupling constants as the initial conditions. We first introduce weak coupling between the ladder and the bulk into each phase of the ladder, and plot the solutions in Fig. 4. We find that the consequent RG flow and the stability of each phase are fully consistent with the above perturbative analysis. However, as shown in Fig. 5, a stronger coupling to the bulk can destabilize the singlet phase of the ladder and drive $K_{\mathrm{S}}$, $K_{\mathrm{A}}$, $g_{1}$ and $g_{xx}^{-}$ to diverge and other coupling constants to vanish, which turns out to be an AF state. Therefore, increasing the coupling between the ladder and the bulk state can induce a direct transition from the ordinary phase to the extraordinary phase with AF order on the boundary, which is consistent with the boundary transition of the dangling ladder system observed in the numerical simulations [30].
IV Irrelevance of weak velocity difference